OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (0, 1/2) and (0, 3/4). is it (0, 2/3)?

OpenStudy (lgbasallote):

how did you get 2/3?

OpenStudy (lgbasallote):

did you do \[\huge \\frac 12 \div \frac 34?\]

OpenStudy (lgbasallote):

i mean \[\frac 12 \div \frac 34\]

OpenStudy (anonymous):

(0+0) / 2 = 0 1/2 + 3/4 = 2/4 + 3/4 = 5/4 5/4 / 2 = 2.5/4 or 5/8 answer should be easy to find now

OpenStudy (anonymous):

(0, 5/8)

OpenStudy (anonymous):

@lgbasallote is that 1/24?